It was experimentally determined from three trials that an object had a density of 0.96, 0.99, and 0.95 g/mL. What is the average density and standard deviation of this object?

a. 0.967 (±0.021) g/mL

b. 0.967 (±0.02) g/mL

c. 0.97 (±0.02) g/mL

d. 0.96 (±0.02) g/mL
Answer #1

Average density (av) = sum of all measurements / number of measurements = (0.96 + 0.99 + 0.95) / 3 = 0.96

Standard deviation (SD) = [(av - measurement)2 / number of measurements]1/2

= [{(0.96 - 0.96)2 + (0.96 - 0.99)2 + (0.96 - 0.95)2 } / 3]1/2 = 0.02 g/ml

Answer is d. 0.96 (+- 0.02) g/ml